OpenStudy (anonymous):

The value of 26- 16 / 2 + 3 * 4 =

OpenStudy (anonymous):

26- 16 / 2 + 3 * 4 = 26-8+12=38-8=30

OpenStudy (anonymous):

@jbri555

OpenStudy (anonymous):

thank you @dpasingh

OpenStudy (anonymous):

Need help ?